Find the momenta of the following problems.

a 20-kg bicycle moving at 7m/s northward


1
Expert's answer
2021-11-15T14:00:40-0500
"p=mv=20\ kg\times7\ \dfrac{m}{s}=140\ \dfrac{kg\times m}{s}."